40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Mar 22 16:29:04 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
    Reply  Thu Mar 24 19:07:38 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
       Reply  Sat Mar 26 13:36:09 2016, tara, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT IMG_2468.JPGIMG_2466.JPG
          Reply  Tue Mar 29 23:23:37 2016, rana, DailyProgress, PMC, Epoxy glue on PMC mirror and PZT 
             Reply  Wed Mar 30 08:29:15 2016, tara, DailyProgress, PMC, test on new PMC IMG_2528.JPGIMG_2540.JPG
                Reply  Wed Mar 30 10:28:44 2016, Evan, DailyProgress, PMC, test on new PMC 
                   Reply  Thu Mar 31 15:11:27 2016, tara, DailyProgress, PMC, test on new PMC 
                      Reply  Tue Apr 5 16:32:36 2016, tara, DailyProgress, PMC, test on new PMC 
                         Reply  Wed Apr 6 15:02:50 2016, tara, DailyProgress, PMC, test on new PMC 
                            Reply  Fri Apr 8 09:44:49 2016, tara, DailyProgress, PMC, test on new PMC IMG_2634.JPGIMG_2635.JPGIMG_2633.JPG
Message ID: 1627     Entry time: Fri Apr 8 09:44:49 2016     In reply to: 1624
Author: tara 
Type: DailyProgress 
Category: PMC 
Subject: test on new PMC 

I checked the VME crate, and made sure that the PMC cards/ LO were functioning.

There was a problem with the VME crate and we couldn't turn on the kepco power supply for +/- 24 V for the PMC servo cards and 21.5 MHz LO on the crate. There were wires that connected to the power supply and the loose ends shorted together. I fixed that by taping over the wires. Now the power is up +/- 24V.

Summary of the crate status:

  • the crate is powered up with +/-24V. The current from each power supply is ~0.8 A. 
  • The power cables to TTFSS in the crate (+/- 17 and 24 V) are disconnected.
  • Both PMC servo cards, are working. Though we cannot adjust the gain/ switches.
  • The 21.5 MHz LO is working. But without the computer, we cannot adjust the power and phase (currently ~ -35 dBm). The signal is not very sinusoidal. (I think I reported this before)

 

Quote:
  • Note: V input for RFPD should be +/- 15 V (schematic)
  • Power for driving 21.5 MHz resonant EOM, can be up to 20 dBm , PSL:791 (seems that we have used only 9 dBm from the LO driver card, PSL:1092 )
  • power for mixer (LO) , 23 dBm ( Minicircuit Ray-3, see the schematic)
  • Schematic for LO for PMC , DCC

 

ELOG V3.1.3-